# The Mixed Integer Trust Region Problem

### v1

@@ 3.2 Far points
$=(v+g)^{\mathsf{T}}H(v+g)+h^{\mathsf{T}}(v+g)$
$=v^{\mathsf{T}}Hv+2v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+h^{\mathsf{T}}v+h^{\mathsf{T}}g$
$\geq v^{\mathsf{T}}Hv+2v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+h^{\mathsf{T}}g$
$\geq v^{\mathsf{T}}Hv-\lvert 2v^{\mathsf{T}}Hg\rvert-\lvert g^{\mathsf{T}}Hg\rvert-\lvert h^{\mathsf{T}}g\rvert$
$\geq(1-\epsilon)\lVert H\rVert_{2}-2\lVert v\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}-\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}-\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\geq(1-\epsilon)\lVert H\rVert_{2}-2\epsilon\lVert H\rVert_{2}-\epsilon^{2}\lVert H\rVert_{2}-\epsilon\lVert H\rVert_{2}$
$=\left(1-4\epsilon-\epsilon^{2}\right)\lVert H\rVert_{2}.$
@@
$f(w)-f(u)$ $\geq\left(1-9\epsilon/2-5\epsilon^{2}/4\right)\lVert H\rVert_{2}.$
@@
$\geq-\epsilon^{2}\lVert H\rVert_{2}/4-\epsilon\lVert H\rVert_{2}/2$
$=-\left(\epsilon/2+\epsilon^{2}/4\right)\lVert H\rVert_{2}.$
Here, the third inequality holds because $\lVert u\rVert_{2}\leq\epsilon/2$ and $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. Next, we give an upper bound for $f(w)$. Let $g\in\mathbb{R}^{n}$ so that $w=v+g$. Since $w\in\mathcal{B}(v,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
@@
$=(v+g)^{\mathsf{T}}H(v+g)+h^{\mathsf{T}}(v+g)$
$=v^{\mathsf{T}}Hv+2v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+h^{\mathsf{T}}v+h^{\mathsf{T}}g$
$\leq v^{\mathsf{T}}Hv+2v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+h^{\mathsf{T}}g$
$\leq v^{\mathsf{T}}Hv+\lvert 2v^{\mathsf{T}}Hg\rvert+\lvert g^{\mathsf{T}}Hg\rvert+\lvert h^{\mathsf{T}}g\rvert$
$\leq-(1-\epsilon)\lVert H\rVert_{2}+2\lVert v\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}+\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}+\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\leq-(1-\epsilon)\lVert H\rVert_{2}+2\epsilon\lVert H\rVert_{2}+\epsilon^{2}\lVert H\rVert_{2}+\epsilon\lVert H\rVert_{2}$
$=-\left(1-4\epsilon-\epsilon^{2}\right)\lVert H\rVert_{2}.$
@@
$f(u)-f(w)$ $\geq\left(1-9\epsilon/2-5\epsilon^{2}/4\right)\lVert H\rVert_{2}.$
4 Far points for Equation MITR
In this section we study Equation MITR. The main goal is to introduce an algorithm that either finds two feasible solutions whose distance in objective value is close to $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$, or finds a nonzero vector $d$ along which the feasible region is flat. This algorithm will be a key component in the design of the approximation algorithms for Equations E-MIQP and MIQP stated in Theorems 1 and 2. To state this result, we need to introduce the concept of width. Let $\mathcal{S}\subseteq\mathbb{R}^{n}$ be a bounded closed set and let $d\in\mathbb{R}^{n}$. We define the width of $\mathcal{S}$ along $d$ to be
@@ 4 Far points for Equation MITR
We are now ready to state the main result of this section.
Proposition 1 (Far points for Equation MITR).
There is an algorithm which, for any $\epsilon\in(0,1/4]$, either finds two feasible solutions $x^{1},x^{2}$ to Equation MITR such that
$f(x^{1})-f(x^{2})\geq\left(1-9\epsilon/2-5\epsilon^{2}/4\right)\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\},$
or finds a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}(0,1))\leq 2p2^{p(p-1)/4}/\epsilon$. The running time of the algorithm is polynomial in the size of Equation MITR and in $\log(1/\epsilon)$.
@@ 4.1 Ingredients
Let $a\in\mathbb{Q}$ be positive. There is an algorithm which, for any positive $\epsilon\in\mathbb{Q}$, finds a positive $l\in\mathbb{Q}$ such that $l\leq\sqrt{a}\leq(1+\epsilon)l$. The running time of the algorithm is polynomial in $\operatorname{size}(a),\log(1/\epsilon)$.
@@
We use Lemma 1 with $k:=\lceil\log(1/\epsilon)+3/2\operatorname{size}(a)+1\rceil$. We only need to show that the obtained $l,u$ satisfy $l>0$ and $u\leq(1+\epsilon)l$.
Using $k\geq 3/2\operatorname{size}(a)+1$ and $2^{-\operatorname{size}(a)}\leq a\leq 2^{\operatorname{size}(a)}$, we obtain
$\begin{split}l&=u-\frac{\max\left\{1,a\right\}}{2^{k}}\geq\sqrt{a}-\frac{\max\left\{1,a\right\}}{2^{3/2\operatorname{size}(a)+1}}\geq 2^{-\operatorname{size}(a)/2}-\frac{2^{\operatorname{size}(a)}}{2^{3/2\operatorname{size}(a)+1}}\\ &=2^{-\operatorname{size}(a)/2}-2^{-\operatorname{size}(a)/2-1}=2^{-\operatorname{size}(a)/2-1},\end{split}$ (1)
@@
We now show $u\leq(1+\epsilon)l$. Using $k\geq\log(1/\epsilon)+3/2\operatorname{size}(a)+1$, $a\leq 2^{\operatorname{size}(a)}$, and (1) we have
$u$ $=l+\frac{\max\left\{1,a\right\}}{2^{k}}\leq l+\frac{2^{\operatorname{size}(a)}}{2^{\log(1/\epsilon)+3/2\operatorname{size}(a)+1}}$
$=l+2^{-\operatorname{size}(a)/2-1}2^{\log(\epsilon)}\leq l+l\epsilon=(1+\epsilon)l.$
The second ingredient is a Lenstra-type result, which follows with little effort from proposition 4 in [5].
@@
In the first case, the vector $\bar{x}^{\prime}:=\bar{x}+c$ is in $\mathcal{B}(a,\delta)\cap(\Lambda+\operatorname{span}(\Lambda)^{\perp}+\{c\})$, and we have $\Lambda+\operatorname{span}(\Lambda)^{\perp}=\Pi_{p}(b^{1},\dots,b^{n})$. In the second case, $d\in\operatorname{span}(\Lambda)$ implies $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$. We then have $d^{\mathsf{T}}c^{i}=d^{\mathsf{T}}b^{i}$, for $i=1,\dots,p$, thus $d^{\mathsf{T}}c^{1},\dots,d^{\mathsf{T}}c^{p}$ integer implies $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer. Furthermore, by definition of width, we obtain $\operatorname{width}_{d}(\mathcal{B}(a,\delta))=\operatorname{width}_{d}(\mathcal{B}(a-c,\delta))\leq p2^{p(p-1)/4}$. ∎
@@ 4.2 Proof of Proposition 1
We are now ready to show Proposition 1. Before presenting its proof, we briefly discuss the the idea of the algorithm. We consider separately the cases $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$ and $\lVert h\rVert_{2}>\lVert H\rVert_{2}$. In the case $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$, we would like to consider a vector $v\in\mathcal{B}(0,1)$ with $|v^{\mathsf{T}}Hv|=\lVert H\rVert_{2}$. However, this vector could be irrational, thus we find instead a vector $v\in\mathcal{B}(0,1)$ with $|v^{\mathsf{T}}Hv|$ that is $\epsilon$-close to $\lVert H\rVert_{2}$. Next, we use our flatness lemma to seek two feasible points $u,w$ to Equation MITR that are $\epsilon$-close to the origin and to $v$, respectively. If we find both these vectors, we show that they have a difference in objective value close to $\lVert H\rVert_{2}$. Otherwise, we show that we have found a nonzero vector ($d$ in the statement) along which the feasible region $\mathcal{B}(0,1)$ is flat.
In the remaining case $\lVert h\rVert_{2}>\lVert H\rVert_{2}$, we would like to consider the vector $h/\lVert h\rVert_{2}$. However, $\lVert h\rVert_{2}$ could be irrational, thus we find instead a scalar $l$ such that $lh\in\mathcal{B}(0,1)$ is $\epsilon$-close to $h/\lVert h\rVert_{2}$. Next, we use our flatness lemma to seek two feasible points $s,t$ to Equation MITR that are $\epsilon$-close to $lh$ and to $-lh$, respectively. If we find both these vectors, we show that they have a difference in objective value close to $\lVert h\rVert_{2}$. Otherwise, we show that we have found a nonzero vector ($d$ in the statement) along which the feasible region $\mathcal{B}(0,1)$ is flat. We now give the complete proof of Proposition 1, which uses Lemmas 8, 5, 7, 6 and 9.
@@
We start with the case $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$. Let $\mathcal{B}_{u}:=\mathcal{B}(0,\epsilon/2)$. From Lemma 9, we either find a vector $u\in\mathcal{B}_{u}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{u})\leq p2^{p(p-1)/4}$. Using the algorithm in Lemma 5, we find a vector $v\in\mathcal{B}(0,1)$ such that $|v^{\mathsf{T}}Hv|\geq(1-\epsilon)\lVert H\rVert_{2}$. Note that the vector $-v$ also lies in $\mathcal{B}(0,1)$ and we have $(-v)^{\mathsf{T}}H(-v)=v^{\mathsf{T}}Hv$. Hence, by eventually redefining $v:=-v$, we now assume $h^{\mathsf{T}}v\geq 0$ if $v^{\mathsf{T}}Hv\geq 0$, and $h^{\mathsf{T}}v\leq 0$ if $v^{\mathsf{T}}Hv<0$. Let $\mathcal{B}_{w}:=\mathcal{B}((1-\epsilon/2)v,\epsilon/2)$. From Lemma 9, we either find a vector $w\in\mathcal{B}_{w}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{w})\leq p2^{p(p-1)/4}$.
Consider the subcase where Lemma 9 successfully found both vectors $u$ and $w$. Note that $\mathcal{B}_{u}$ and $\mathcal{B}_{w}$ are contained in $\mathcal{B}(0,1)$, thus $u,w$ are both feasible to Equation MITR. We also have $\mathcal{B}_{w}\subseteq\mathcal{B}(v,\epsilon)$, thus we can apply Lemma 7 and obtain
$\lvert f(w)-f(u)\rvert\geq\left(1-9\epsilon/2-5\epsilon^{2}/4\right)\lVert H\rVert_{2}.$
Next, consider the subcase where Lemma 9 did not find both vectors $u$ and $w$. In this case, Lemma 9 found a vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B})\leq p2^{p(p-1)/4}$, where $\mathcal{B}\in\{\mathcal{B}_{u},\mathcal{B}_{w}\}$. We then have
$\operatorname{width}_{d}(\mathcal{B}(0,1))$ $=\frac{2}{\epsilon}\operatorname{width}_{d}(\mathcal{B})\leq\frac{2}{\epsilon}p2^{p(p-1)/4}.$

### v2

@@ 4.2 Feasible points to Equation TR with distant objective values
$=((1-\epsilon)v+g)^{\mathsf{T}}H((1-\epsilon)v+g)+h^{\mathsf{T}}((1-\epsilon)v+g)$
$=(1-\epsilon)^{2}v^{\mathsf{T}}Hv+2(1-\epsilon)v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+(1-\epsilon)h^{\mathsf{T}}v+h^{\mathsf{T}}g$
$\geq(1-\epsilon)^{2}v^{\mathsf{T}}Hv+2(1-\epsilon)v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+h^{\mathsf{T}}g$
$\geq(1-\epsilon)^{2}v^{\mathsf{T}}Hv-\lvert 2(1-\epsilon)v^{\mathsf{T}}Hg\rvert-\lvert g^{\mathsf{T}}Hg\rvert-\lvert h^{\mathsf{T}}g\rvert$
$\geq(1-\epsilon)^{3}\lVert H\rVert_{2}-2(1-\epsilon)\lVert v\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}-\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}-\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\geq(1-\epsilon)^{3}\lVert H\rVert_{2}-2\epsilon(1-\epsilon)\lVert H\rVert_{2}-\epsilon^{2}\lVert H\rVert_{2}-\epsilon\lVert H\rVert_{2}$
$=\left(1-6\epsilon+4\epsilon^{2}-\epsilon^{3}\right)\lVert H\rVert_{2}.$
@@
$f(w)-f(u)$ $\geq\left(1-7\epsilon+3\epsilon^{2}-\epsilon^{3}\right)\lVert H\rVert_{2}$
$\geq\left(1-7\epsilon\right)\lVert H\rVert_{2}.$
@@
$\geq-\epsilon^{2}\lVert H\rVert_{2}-\epsilon\lVert H\rVert_{2}$
$=-\left(\epsilon+\epsilon^{2}\right)\lVert H\rVert_{2}.$
Here, the third inequality holds because $\lVert u\rVert_{2}\leq\epsilon$ and $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. Next, we give an upper bound for $f(w)$. Let $g\in\mathbb{R}^{n}$ so that $w=(1-\epsilon)v+g$. Since $w\in\mathcal{B}((1-\epsilon)v,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
@@
$=((1-\epsilon)v+g)^{\mathsf{T}}H((1-\epsilon)v+g)+h^{\mathsf{T}}((1-\epsilon)v+g)$
$=(1-\epsilon)^{2}v^{\mathsf{T}}Hv+2(1-\epsilon)v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+(1-\epsilon)h^{\mathsf{T}}v+h^{\mathsf{T}}g$
$\leq(1-\epsilon)^{2}v^{\mathsf{T}}Hv+2(1-\epsilon)v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+h^{\mathsf{T}}g$
$\leq(1-\epsilon)^{2}v^{\mathsf{T}}Hv+\lvert 2(1-\epsilon)v^{\mathsf{T}}Hg\rvert+\lvert g^{\mathsf{T}}Hg\rvert+\lvert h^{\mathsf{T}}g\rvert$
$\leq-(1-\epsilon)^{3}\lVert H\rVert_{2}+2(1-\epsilon)\lVert v\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}+\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}+\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\leq-(1-\epsilon)^{3}\lVert H\rVert_{2}+2\epsilon(1-\epsilon)\lVert H\rVert_{2}+\epsilon^{2}\lVert H\rVert_{2}+\epsilon\lVert H\rVert_{2}$
$=-\left(1-6\epsilon+4\epsilon^{2}-\epsilon^{3}\right)\lVert H\rVert_{2}.$
@@
$f(u)-f(w)$ $\geq\left(1-7\epsilon+3\epsilon^{2}-\epsilon^{3}\right)\lVert H\rVert_{2}$
$\geq\left(1-7\epsilon\right)\lVert H\rVert_{2}.$
5 Feasible points to Equation MITR with distant objective values
In this section we study Equation MITR. The main goal is to introduce an algorithm that either finds two feasible solutions whose difference in objective value is at least $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$, for $\epsilon$ that goes to zero, or finds a nonzero vector $d$ along which the feasible region is flat. This algorithm will be a key component in the design of the approximation algorithms for Equations E-MIQP and MIQP stated in Theorems 1 and 2. The reason we seek a bound featuring $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$ will become clear later, in the proof of Proposition 2, and is essentially because the same $\max$ appears in the bound in Lemma 4. To state the main result of this section, we need to introduce the concept of width. Let $\mathcal{S}\subseteq\mathbb{R}^{n}$ be a nonempty bounded closed set and let $d\in\mathbb{R}^{n}$. We define the width of $\mathcal{S}$ along $d$ to be
@@ 5 Feasible points to Equation MITR with distant objective values
Proposition 1 (Distant points for Equation MITR).
There is an algorithm which, for any $\epsilon\in(0,1/3]$, either finds two feasible solutions $x^{1},x^{2}$ to Equation MITR such that
$f(x^{1})-f(x^{2})\geq\left(1-7\epsilon\right)\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\},$
or finds a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}(0,1))\leq p2^{p(p-1)/4}/\epsilon$. The running time of the algorithm is polynomial in the size of Equation MITR and in $\log(1/\epsilon)$.
@@ 5.1 Ingredients
Let $\alpha\in\mathbb{Q}$ be positive. There is an algorithm which, for any positive $\epsilon\in\mathbb{Q}$, finds a positive $l\in\mathbb{Q}$ such that $l\leq\sqrt{\alpha}\leq(1+\epsilon)l$. The running time of the algorithm is polynomial in $\operatorname{size}(\alpha),\log(1/\epsilon)$.
@@
We use Lemma 1 with $k:=\lceil\log(1/\epsilon)+3\operatorname{size}(\alpha)/2+1\rceil$. We only need to show that the obtained $l,u$ satisfy $l>0$ and $u\leq(1+\epsilon)l$.
Using $k\geq 3\operatorname{size}(\alpha)/2+1$ and $2^{-\operatorname{size}(\alpha)}\leq\alpha\leq 2^{\operatorname{size}(\alpha)}$, we obtain
$\begin{split}l&=u-\frac{\max\left\{1,\alpha\right\}}{2^{k}}\geq\sqrt{\alpha}-\frac{\max\left\{1,\alpha\right\}}{2^{3\operatorname{size}(\alpha)/2+1}}\geq 2^{-\operatorname{size}(\alpha)/2}-\frac{2^{\operatorname{size}(\alpha)}}{2^{3\operatorname{size}(\alpha)/2+1}}\\ &=2^{-\operatorname{size}(\alpha)/2}-2^{-\operatorname{size}(\alpha)/2-1}=2^{-\operatorname{size}(\alpha)/2-1},\end{split}$ (3)
@@
We now show $u\leq(1+\epsilon)l$. Using $k\geq\log(1/\epsilon)+3\operatorname{size}(\alpha)/2+1$, $\alpha\leq 2^{\operatorname{size}(\alpha)}$, and (3) we have
$u$ $=l+\frac{\max\left\{1,\alpha\right\}}{2^{k}}\leq l+\frac{2^{\operatorname{size}(\alpha)}}{2^{\log(1/\epsilon)+3\operatorname{size}(\alpha)/2+1}}$
$=l+2^{-\operatorname{size}(\alpha)/2-1}2^{\log(\epsilon)}\leq l+l\epsilon=(1+\epsilon)l.$
The second ingredient is a flatness result for balls, which follows with little effort from proposition 4 in [7].
@@
In the first case, the vector $\bar{x}^{\prime}:=\bar{x}+c$ is in $\mathcal{B}(a,\delta)\cap(\Lambda+\operatorname{span}(\Lambda)^{\perp}+\{c\})$, and we have $\Lambda+\operatorname{span}(\Lambda)^{\perp}=\Lambda+\operatorname{span}(b^{p+1},\dots,b^{n})=\Pi_{p}(b^{1},\dots,b^{n})$. In the second case, $d\in\operatorname{span}(\Lambda)$ implies $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$. We then have $d^{\mathsf{T}}c^{i}=d^{\mathsf{T}}b^{i}$, for $i=1,\dots,p$, thus $d^{\mathsf{T}}c^{1},\dots,d^{\mathsf{T}}c^{p}$ integer implies $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer. Furthermore, by definition of width, we obtain $\operatorname{width}_{d}(\mathcal{B}(a,\delta))=\operatorname{width}_{d}(\mathcal{B}(a-c,\delta))\leq p2^{p(p-1)/4}$. ∎
@@ 5.2 Proof of Proposition 1
We are now ready to show Proposition 1. Before presenting its proof, we briefly discuss the the idea of the algorithm. The goal is to either find two feasible solutions to Equation MITR whose difference in objective value is at least $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$, for $\epsilon$ that goes to zero, or to find a nonzero vector $d$ along which the feasible region is flat. Depending on which one among $\lVert H\rVert_{2}$ and $\lVert h\rVert_{2}$ is largest, we search for the two feasible solutions in different regions in $\mathcal{B}(0,1)$; As a result, we consider separately the cases $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$ and $\lVert h\rVert_{2}>\lVert H\rVert_{2}$.
In the case $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$, we would like to consider a vector $v\in\mathcal{B}(0,1)$ with $|v^{\mathsf{T}}Hv|=\lVert H\rVert_{2}$. However, this vector could be irrational, thus we use Lemma 5 to find, instead, a vector $v\in\mathcal{B}(0,1)$ with $|v^{\mathsf{T}}Hv|$ that is $\epsilon$-close to $\lVert H\rVert_{2}$. Next, we use our flatness lemma to seek two feasible points $u,w$ to Equation MITR that are $\epsilon$-close to the origin and to $v$, respectively. If we find both these vectors, we use Lemma 7 to show that they have a difference in objective value that is at least $\lVert H\rVert_{2}$, for $\epsilon$ that goes to zero. Otherwise, we show that we have found a nonzero vector ($d$ in the statement) along which the feasible region $\mathcal{B}(0,1)$ is flat.
In the remaining case $\lVert h\rVert_{2}>\lVert H\rVert_{2}$, we would like to consider the vector $h/\lVert h\rVert_{2}$. However, $\lVert h\rVert_{2}$ could be irrational, thus we find instead a scalar $l$ such that $lh\in\mathcal{B}(0,1)$ is $\epsilon$-close to $h/\lVert h\rVert_{2}$. Next, we use our flatness lemma to seek two feasible points $s,t$ to Equation MITR that are $\epsilon$-close to $lh$ and to $-lh$, respectively. If we find both these vectors, we use Lemma 6 to show that they have a difference in objective value that is at least $\lVert h\rVert_{2}$, for $\epsilon$ that goes to zero. Otherwise, we show that we have found a nonzero vector ($d$ in the statement) along which the feasible region $\mathcal{B}(0,1)$ is flat. We now give the complete proof of Proposition 1, which uses Lemmas 8, 5, 7, 6 and 9.
@@
We start with the case $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$. Let $\mathcal{B}_{u}:=\mathcal{B}(0,\epsilon)$. From Lemma 9 (with $a:=0$ and $\delta:=\epsilon$), we either find a vector $u\in\mathcal{B}_{u}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{u})\leq p2^{p(p-1)/4}$. Using the algorithm in Lemma 5, we find a vector $v\in\mathcal{B}(0,1)$ such that $|v^{\mathsf{T}}Hv|\geq(1-\epsilon)\lVert H\rVert_{2}$. Note that the vector $-v$ also lies in $\mathcal{B}(0,1)$ and we have $(-v)^{\mathsf{T}}H(-v)=v^{\mathsf{T}}Hv$. Hence, by possibly redefining $v:=-v$, we now assume $h^{\mathsf{T}}v\geq 0$ if $v^{\mathsf{T}}Hv\geq 0$, and $h^{\mathsf{T}}v\leq 0$ if $v^{\mathsf{T}}Hv<0$. Let $\mathcal{B}_{w}:=\mathcal{B}((1-\epsilon)v,\epsilon)$. From Lemma 9, we either find a vector $w\in\mathcal{B}_{w}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{w})\leq p2^{p(p-1)/4}$.
Consider the subcase where Lemma 9 successfully found both vectors $u$ and $w$. Note that $\mathcal{B}_{u}$ and $\mathcal{B}_{w}$ are contained in $\mathcal{B}(0,1)$, thus $u,w$ are both feasible to Equation MITR. We apply Lemma 7 and obtain
$\lvert f(w)-f(u)\rvert\geq\left(1-7\epsilon\right)\lVert H\rVert_{2}.$
Next, consider the subcase where Lemma 9 did not find both vectors $u$ and $w$. In this case, Lemma 9 found a vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B})\leq p2^{p(p-1)/4}$, for $\mathcal{B}\in\{\mathcal{B}_{u},\mathcal{B}_{w}\}$. We then have
$\operatorname{width}_{d}(\mathcal{B}(0,1))$ $=\frac{1}{\epsilon}\operatorname{width}_{d}(\mathcal{B})\leq\frac{1}{\epsilon}p2^{p(p-1)/4}.$